### v2

# FreeGraftor: Training-Free Cross-Image Feature Grafting for Subject-Driven Text-to-Image Generation

Abstract.
Subject-driven image generation aims to synthesize novel scenes that faithfully preserve subject identity from reference images while adhering to textual guidance, yet existing methods struggle with a critical trade-off between fidelity and efficiency. Tuning-based approaches rely on time-consuming and resource-intensive subject-specific optimization, while zero-shot methods fail to maintain adequate subject consistency. In this work, we propose FreeGraftor, a training-free framework that addresses these limitations through cross-image feature grafting. Specifically, FreeGraftor employs semantic matching and position-constrained attention fusion to transfer visual details from reference subjects to the generated image. Additionally, our framework incorporates a novel noise initialization strategy to preserve geometry priors of reference subjects for robust feature matching. Extensive qualitative and quantitative experiments demonstrate that our method enables precise subject identity transfer while maintaining text-aligned scene synthesis. Without requiring model fine-tuning or additional training, FreeGraftor significantly outperforms existing zero-shot and training-free approaches in both subject fidelity and text alignment. Furthermore, our framework can seamlessly extend to multi-subject generation, making it practical for real-world deployment. Our code is available at https://github.com/Nihukat/FreeGraftor.
copyright: noneconference: ; ;
Figure 1. Subject-driven generation results of our FreeGraftor, achieved without any training or tuning process. These results demonstrate the superior properties of FreeGraftor, including (a) pixel-level detail preservation, (b) flexible text-guided control, and (c) support for multiple reference subjects.
1. Introduction
The democratization of text-to-image generation has ignited widespread interest in personalized content creation, where users seek to synthesize novel scenes containing specific visual concepts—be it a cherished pet, a unique artifact, or a fictional character. While diffusion models(Nichol et al., 2021; Saharia et al., 2022; Ramesh et al., 2022; Rombach et al., 2022; Podell et al., 2023) excel at generating diverse imagery from textual prompts, adapting them to preserve subject identity without costly per-subject optimization remains a critical challenge.
Current solutions for subject-driven generation(Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023; Li et al., 2023; Ye et al., 2023) face a critical dilemma between fidelity and efficiency. Tuning-based methods, such as DreamBooth(Ruiz et al., 2023) and Textual Inversion(Gal et al., 2022), capture subject identity by fine-tuning models or learning embeddings on user-provided subject datasets. These methods rely on subject-specific optimization to achieve high subject consistency, which are resource-intensive and time-consuming, prone to catastrophic forgetting, and struggle with multi-subject composition. Zero-shot methods, such as IP-Adapter(Ye et al., 2023) and BLIP-Diffusion(Li et al., 2023), extract subject identities from reference images via additional trainable modules (e.g., image encoders or adapters) and integrate them into the text-to-image generation process. While these methods directly generate images containing reference subjects, they often fail to retain fine visual details, especially for novel concepts, as the limited parameter capacity of auxiliary modules fundamentally restricts their ability to reconstruct high-fidelity visual details.
It is worth noting that pre-trained text-to-image models inherently possess powerful visual representation capabilities, as evidenced by their ability to synthesize intricate textures and structures from textual prompts. However, current approaches fail to fully exploit this potential. Tuning-based methods modify the image generation distribution through fine-tuning, while zero-shot approaches offload visual representation learning to auxiliary modules. In contrast, directly tapping into the base model’s own feature space for cross-image knowledge transfer offers a promising way to bridge the gap between fidelity and efficiency. Recent advances in appearance transfer(Epstein et al., 2023; Alaluf et al., 2024; Mou et al., 2023, 2024; Go et al., 2024) lend credence to this idea, demonstrating that diffusion models can transplant visual attributes such as style and texture between images through semantic feature matching. For instance, Eye-for-an-eye(Go et al., 2024) successfully transfers localized appearances by establishing cross-image correspondences in U-Net’s feature space. By computing pixel-wise similarity (detailed in Sec. 3.2), we demonstrate that such cross-image semantic correspondences can also be identified in the feature space of modern MM-DiT-based(Esser et al., 2024) models (e.g., FLUX.1(Labs, 2024)). As illustrated in Figure 2, two pixels that are similar in FLUX.1’s feature space are also semantically corresponding.
Figure 2. Cross-image semantic correspondence in FLUX.1. For each specified pixel in the reference image, we identify the most similar pixel in the target image within FLUX.1‘s feature space. These corresponding pixel pairs are then visually connected by colored lines.
Building on this insight, we propose FreeGraftor, a training-free framework that resolves the fidelity-efficiency dilemma in subject-driven generation via cross-image feature grafting. Our approach extracts visual features from reference images using rectified flow inversion(Rout et al., 2024; Wang et al., 2024b; Deng et al., 2024) and injects them into the generation process. Specifically, we introduce a novel cross-image feature grafting mechanism that transfers appearance details through semantic matching and position-constrained attention fusion in the feature space of text-to-image diffusion models. To bind reference patches to their corresponding generated patches, we implement a position embedding replication strategy, enabling semantically matched patches to share positional information during attention computation. Furthermore, we identify that proper noise initialization is crucial for robust semantic matching and structural preservation of reference subjects. Thus, we propose a structure-consistent initialization strategy: first constructing a collage containing the reference subject, then inverting it as the initialization noise for generation. Our method achieves pixel-level detail preservation (e.g., text, patterns) with fast inference and low computational overhead, requiring no additional training or optimization.
Extensive experiments demonstrate that our method outperforms existing approaches in both subject consistency and text alignment while maintaining competitive efficiency. When combining multiple subjects, FreeGraftor preserves individual identities and harmoniously integrates all subjects into a coherent scene without layout conflicts or attribute confusion.
Our contributions are summarized as follows:
  • We introduce FreeGraftor, a plug-and-play framework for subject-driven image generation that achieves pixel-level detail preservation and flexible text-guided control without fine-tuning or training.
  • We develop a cross-image feature grafting technique that transfers visual characteristics of reference subjects to semantically corresponding regions in generated images via semantic matching and position-constrained attention fusion.
  • We propose a novel noise initialization strategy that preserves reference geometry through automated collage construction and rectified flow inversion, seamlessly extending our method to multi-subject scenarios.
2. Related Work
2.1. Multimodal-Diffusion Transformer
Text-to-image diffusion models have achieved remarkable success in generating high-quality and diverse images, exemplified by DALL-E 2(Ramesh et al., 2022), Imagen(Saharia et al., 2022), GLIDE(Nichol et al., 2021), and eDiff-I(Balaji et al., 2022). Early popular open-source models, such as Stable Diffusion(Rombach et al., 2022) and SDXL(Podell et al., 2023), are based on U-Net(Ronneberger et al., 2015) architecture and utilize cross-attention between image and text to introduce textual conditions. In contrast, Diffusion Transformer (DiT)(Peebles and Xie, 2023) replaces the U-Net backbone with a Transformer(Vaswani et al., 2017), demonstrating superior scalability and performance. Recent advanced open-source models, including Stable Diffusion 3(Esser et al., 2024) and FLUX.1(Labs, 2024), follow this design and project text and image tokens into a unified space to compute joint attention, thereby better integrating complex relationships between text and image modalities. This architecture is termed Multimodal-Diffusion Transformer (MM-DiT). In this work, we implement our method on FLUX.1, where joint attention is computed as:
(1) $Q=\left[Q^{\text{txt}};Q^{\text{img}}\right],\quad K=\left[K^{\text{txt}};K^{\text{img}}\right],\quad V=\left[V^{\text{txt}};V^{\text{img}}\right]$
(2) $\tilde{Q}=PE\odot Q,\quad\tilde{K}=PE\odot K$
(3) $A=\operatorname{softmax}\left(\frac{\tilde{Q}\tilde{K}^{\top}}{\sqrt{d}}\right),\quad H_{\text{out}}=AV$
Here, [;][;][ ; ] denotes concatenation, $Q$, $K$, $V$ represent queries, keys, values, $PE$ is position embedding, $d$ denotes the dimension of the input features, $A$ denotes attention scores, and $H_{\text{out}}$ is the output of the attention layer.
2.2. Subject-Driven Generation
Subject-driven generation(Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023; Li et al., 2023; Ye et al., 2023) aims to synthesize novel scenes containing user-specified subjects. Existing approaches fall into two categories: tuning-based and zero-shot. Tuning-based methods(Dong et al., 2022; Ruiz et al., 2023; Gal et al., 2022; Voynov et al., 2023; Alaluf et al., 2023; Liu et al., 2023; Kumari et al., 2023; Gu et al., 2023) achieve subject-driven generation by customizing models, typically requiring per-subject fine-tuning(Ruiz et al., 2023; Kumari et al., 2023; Gu et al., 2023) or iterative learning of unique embeddings(Gal et al., 2022; Voynov et al., 2023; Alaluf et al., 2023; Liu et al., 2023). While subject-specific tuning effectively captures visual details of the reference subject, it is resource-intensive and time-consuming, risks overfitting and catastrophic forgetting, and struggles with multi-subject composition. Zero-shot methods(Gal et al., 2023; Wei et al., 2023; Jia et al., 2023; Shi et al., 2024; Arar et al., 2023; Hua et al., 2023; Li et al., 2023; Ye et al., 2023; Ruiz et al., 2024; Wang et al., 2024a; Xiao et al., 2024) introduce additional learnable modules to process extra image inputs, seamlessly integrating visual information of the reference subject into the generation process. These methods enable personalization at inference time but sacrifice fine visual details, especially for novel concepts. Furthermore, they require large-scale retraining when switching to new base models. Moreover, both categories can degrade the performance of base models, resulting in reduced quality of generated images.
Beyond these categories, several training-free approaches(Ding et al., 2024; Shin et al., 2024) aim to harness the inherent capabilities of pre-trained text-to-image models for plug-and-play subject-driven generation. For example, FreeCustom(Ding et al., 2024) injects reference features by concatenating key and value pairs of reference and generated images in attention layers, emphasizing visual characteristics of the reference subject through attention score editing. However, this approach relies on layout alignment across multiple reference images and suffers from visual distortion due to coarse attention manipulation. DiptychPrompting(Shin et al., 2024) creatively reframes the subject-driven generation task as a diptych inpainting process, utilizing the visual context processing capabilities of pre-trained models to achieve high subject consistency. Unfortunately, pixel-space diptych construction incurs significant memory overhead, limiting its applicability to high-resolution images (e.g., 1024×1024 or larger). Additionally, this paradigm does not support generation with multiple reference subjects. In contrast, our work identifies cross-image semantic correspondences in MM-DiT’s feature space and binds matched patches via position embedding manipulation, enabling efficient subject feature injection with minimal computational cost.
2.3. Appearance Transfer
Appearance transfer methods(Epstein et al., 2023; Alaluf et al., 2024; Mou et al., 2023, 2024; Go et al., 2024) aim to combine the appearance of one image with the structure of another. Cross-Image(Alaluf et al., 2024) inject key-value pairs across images to transfer appearance from reference to target positions in the generated image. Dragon-diffusion(Mou et al., 2023) and DiffEditor(Mou et al., 2024) integrate noise guidance with key-value injection. However, these methods struggle to preserve structure because attention scores do not necessarily reflect semantic correspondences between reference and target images, potentially leading to erroneous matches. DIFT(Tang et al., 2023) identifies fine-grained semantic correspondences in U-Net feature space by computing pixel-pair similarities. Building on this, Eye-for-an-Eye(Go et al., 2024) achieves appearance transfer by directly replacing features of corresponding pixels. Nevertheless, these U-Net-based approaches are incompatible with modern MM-DiT architectures. Moreover, these methods focus on style transfer rather than identity preservation, which will cause geometric inconsistencies (e.g., shape, limb proportions) between reference and generated subjects if directly applied to subject-driven generation. In this work, we explore semantic matching and fine-grained feature transfer in MM-DiT while employing a noise initialization strategy to preserve the geometric structure of reference subjects.
Figure 3. Overview of FreeGraftor. First, we construct a collage based on the given text prompt and the reference subject image (Stage 1). Next, we invert this collage and record its diffusion trajectory (Stage 2). Finally, using the inverted noise as the initial latent representation, our FreeGraftor synthesizes the output image through iterative denoising. During this process, the Semantic-aware Feature Grafting (SAFG) module integrates features from the collage to ensure alignment with the reference subject (Stage 3).
Figure 4. Illustration of the Proposed Semantic-aware Feature Grafting (SAFG) Module. For each patch in the reference image, it first establishes semantic correspondence in the generated image via feature matching. The module then (1) concatenates the key and value pairs of the reference patch with those of the corresponding generated patch, while (2) copying and applying the position embedding from the generated patch to the reference patch’s key. This mechanism enables effective positional information sharing between corresponding patches while maintaining semantic alignment.
3. Method
3.1. Overview
Our framework comprises three stages: collage construction, inversion, and generation (illustrated in Figure 3). First, we construct a collage using the given reference image and text prompt. Next, we invert the collage to obtain initialization noise for the generation phase while extracting reference features. Finally, these features are injected into the generation process to synthesize the final image containing the reference subject within the text-described scene.
3.2. Semantic-Aware Feature Grafting
Most subject-driven generation methods rely on learnable parameters to inject reference features into the generation process. Tuning-based approaches embed subject-specific information into pre-trained models, while zero-shot methods employ additional modules for feature extraction and injection—both incurring significant computational overhead.
Inspired by appearance transfer techniques(Epstein et al., 2023; Alaluf et al., 2024; Mou et al., 2023, 2024; Go et al., 2024), we leverage the inherent capabilities of text-to-image diffusion models to transfer visual characteristics of reference subjects to corresponding regions in generated images without incurring additional training costs. To achieve this, we propose the Semantic-Aware Feature Grafting (SAFG) module, which first establishes semantic correspondences between reference and generated patches via semantic matching, then fuses features through position-constrained attention fusion (Figure 4).
3.2.1. Semantic Matching
Early appearance transfer methods(Alaluf et al., 2024; Mou et al., 2023, 2024) directly concatenate keys and values of reference and generated images to compute cross-image attention. However, high attention scores do not necessarily indicate semantic correspondences across images, and erroneous matches may lead to structural distortions and visual artifacts. DIFT(Tang et al., 2023) revealed that U-Net-based diffusion models contain rich semantic information in their feature space and established cross-image correspondences via cosine similarity-based semantic matching. We extend this principle to MM-DiT-based diffusion models.
Given a reference image $I^{\text{ref}}$ and a predefined binary mask $M^{\text{ref, pre}}$ indicating the reference subject, we extract attention features using rectified flow inversion(Rout et al., 2024; Wang et al., 2024b; Deng et al., 2024) and match reference $H^{\text{ref}}$ and generated $H^{\text{gen}}$ features at the same timestep and attention layer. For each reference patch $i$ within $M^{\text{ref, pre}}$, we compute cosine similarity with all generated patches $j$:
(4) $s_{ij}=\frac{H_{i}^{\text{img,ref}}\cdot H_{j}^{\text{img,gen}}}{\|H_{i}^{\text{img,ref}}\|_{2}\|H_{j}^{\text{img,gen}}\|_{2}},\quad m(i)=\arg\max_{j}s_{ij}$
where $H_{i}^{\text{img,ref}}$ and $H_{j}^{\text{img,gen}}$ denote features of the $i$-th reference patch and $j$-th generated patch at the attention layer input. To ensure reliable matches, we apply two filtering strategies:
i. Similarity Threshold Filtering. For each reference patch $i$, if its maximum cosine similarity $s_{i,m(i)}$ falls below a predefined threshold $\tau$, the match is discarded. This is represented by the binary mask:
(5) $M_{i}^{\text{ref, sim}}=\begin{cases}1,&s_{i,m(i)}\geq\tau\\ 0,&s_{i,m(i)}<\tau\end{cases}$
ii. Cycle Consistency Filtering. Following DreamMatcher(Nam et al., 2024), for each reference patch $i$ (satisfying $M_{i}^{\text{ref, pre}}=1$), we find the most similar reference patch $k$ to its matched generated patch $m(i)$:
(6) $m^{-1}(m(i))=\arg\max_{k:M_{k}^{\text{ref,pre}}=1}\frac{H_{m(i)}^{\text{img,gen}}\cdot H_{k}^{\text{img,ref}}}{\|H_{m(i)}^{\text{img,gen}}\|_{2}\|H_{k}^{\text{img,ref}}\|_{2}}$
Let $\mathbf{p}_{i}^{\text{ref}}$ and $\mathbf{p}_{m^{-1}(m(i))}^{\text{ref}}$ denote the 2D coordinates of patch $i$ and its reverse-matched counterpart. Matches are rejected if their L2 distance exceeds $\delta$:
(7) $d_{i}=\|\mathbf{p}_{i}^{\text{ref}}-\mathbf{p}_{m^{-1}(m(i))}^{\text{ref}}\|_{2},\quad M_{i}^{\text{ref, consi}}=\begin{cases}1,&d_{i}\leq\delta\\ 0,&d_{i}>\delta\end{cases}$
Combining these filters with the predefined mask yields the final reference image mask:
(8) $M^{\text{ref}}=M^{\text{ref, pre}}\odot M^{\text{ref, sim}}\odot M^{\text{ref,consi}}$
3.2.2. Position-Constrained Attention Fusion
After obtaining the filtered reference mask $M^{\text{ref}}$, we transfer features from the retained reference patches. A naive approach is to directly replace the features of the corresponding patches. However, this may introduce artifacts along object boundaries or even cause image distortion (as shown in Figure 7). Unlike U-Net, which models positional relationships through convolutional operations, MM-DiT distinguishes patch proximity and orientation via position embeddings (PE). This enables us to align reference patches with their target positions in the generated image by editing position embeddings.
Specifically, we concatenate both the keys and values of the reference patch to those of the generated patch, along with the position embeddings corresponding to the generated patch associated with the reference patch. This design enables the reference patch and its corresponding generated patch to share identical positional information. Let $K^{\text{txt}}$ and $V^{\text{txt}}$ denote the text embeddings’ keys and values, respectively; $K^{\text{img, gen}}$ and $V^{\text{img, gen}}$ represent the generated image’s keys and values; and $K_{M}^{\text{img, ref}}$ and $V_{M}^{\text{img, ref}}$ correspond to the keys and values of the reference patches satisfying $M^{\text{ref}}$. We concatenate these as follows:
(9) $K^{\text{cat}}=[K^{\text{txt}};K^{\text{img,gen}};K_{M}^{\text{img,ref}}],$
$V^{\text{cat}}=[V^{\text{txt}};V^{\text{img,gen}};V_{M}^{\text{img,ref}}]$
For each retained reference patch $i$, we extract the corresponding position embedding $PE_{m(i)}^{\text{img, gen}}$ based on the matching relationship $m(i)$. The new position embedding is then formed by concatenating:
(10) $PE^{\text{cat}}=[PE^{\text{txt}};PE^{\text{img,gen}};\{PE_{m(i)}^{\text{img,gen}}\mid M_{i}^{\text{ref}}=1\}]$
The revised attention computation is formulated as:
(11) $\tilde{Q}=PE\odot Q,\quad\tilde{K}^{\text{cat}}=PE^{\text{cat}}\odot K^{\text{cat}}$
(12) $A^{+}=\text{softmax}\left(\frac{\tilde{Q}(\tilde{K}^{\text{cat}})^{\top}}{\sqrt{d}}\right),\quad H_{\text{out}}^{+}=A^{+}V^{\text{cat}}$
Here, $A^{+}$ represents the updated attention scores, and $H_{\text{out}}^{+}$ denotes the output of the attention layer. By integrating semantic matching and the sharing of position embeddings, our Semantic-Aware Feature Grafting module effectively binds features between matching patches in the reference and generated images, facilitating cross-image feature grafting.
3.3. Structure-Consistent Initialization
Attention-based feature migration can replicate the appearance of reference subjects (e.g., color, texture) but may not preserve structural details (e.g., shape, body proportions). Additionally, layouts generated from Gaussian noise exhibit significant randomness, leading to discrepancies in shape and size between generated subjects and those in the reference image, potentially causing semantic matching failures. To maintain structural integrity and enhance the robustness of semantic matching, we propose a structure-consistent initialization strategy. This involves creating a collage containing the reference subject and then employing an inversion technique to derive the initial noise required for the generation phase.
As illustrated in Figure 3, given a text prompt (e.g., ”a cat sitting on a chair wearing a birthday hat”), we first use a base text-to-image model to generate a template image that faithfully represents the described scene. To preserve the structure of the reference subject, we replace the corresponding subject in the template image with the reference subject. Specifically, a grounding model localizes the target subject in the template, a segmentation model extracts its mask, and an inpainting model removes the subject from the template. The reference subject is then cropped (using the same grounding and segmentation models), resized, and pasted into the erased region to form a collage. This collage replaces the original reference image as the new reference. Finally, we invert this collage to obtain the initial noise for the generation phase and record the diffusion trajectory during inversion to facilitate feature grafting.
To prevent the generated subject from overfitting the pose of the reference subject, we introduce a dynamic feature dropout strategy. Since structural layouts are predominantly determined during early diffusion steps(Balaji et al., 2022; Alaluf et al., 2024), we reduce reference feature injection in early steps while retaining late-step features to preserve visual details. At timestep $t$, for each reference patch, we apply dropout with probability $p=\omega t$, where $\omega\in[0,1]$ controls dropout intensity. The dropout mask is represented by a Bernoulli distribution:
(13) $M^{\text{drop}}_{i,j}\sim\text{Bernoulli}(1-\omega t)\quad\forall i,j$
The final reference mask is computed by applying dropout to the reference image mask obtained from semantic matching (Sec. 3.2):
(14) ${M^{\text{ref}}}^{\prime}=M^{\text{ref}}\odot M^{\text{drop}}$
This strategy preserves the reference subject’s structure while aligning it with the text-driven layout and maintaining pose flexibility. For multi-subject generation, we iteratively replace subjects in the template and invert the multi-subject collage, avoiding simultaneous processing of multiple reference images and eliminating computational overhead.
Figure 5. Generation results with single reference subject using different methods. Our FreeGraftor achieves pixel-level detail preservation (e.g., text and patterns) while allowing flexible text-guided control (e.g., poses of teddy bears and dogs).
Figure 6. Generation results with multiple reference subjects using different methods. Our FreeGraftor preserves the visual details of all reference subjects while producing high-quality images well-aligned with the text.
Figure 7. Generation results of different ablation variants. (a) Without structure-consistent initialization. (b) Without feature grafting. (c) Using direct key-value concatenation in attention layers instead of semantic matching. (d) Using direct patch feature replacement instead of position-constrained attention fusion. (e) Our full method.
Figure 8. Generation results under different similarity threshold ($\tau$) and cycle-consistency threshold ($\delta$) settings. Higher $\tau$ and lower $\delta$ lead to more stringent filtering.
Figure 9. Generation results under different dropout intensity ($\omega$) settings. Higher $\omega$ improves text alignment but degrades subject consistency.
4. Experiments
4.1. Experimental Setup
4.1.1. Implementation Details
Our method is implemented on FLUX.1-dev. For collage construction, we use Grounding DINO(Liu et al., 2024) to localize subjects, SAM(Kirillov et al., 2023) to segment subjects, and LaMa(Suvorov et al., 2022) to erase the original subject in the template image. For rectified flow inversion, FireFlow(Deng et al., 2024) is employed as the inversion method. Template generation, inversion, and final generation each take 25 diffusion steps. During semantic matching, we set the similarity threshold $\tau=0.2$, cycle consistency threshold $\delta=1.5$, and dropout intensity $\omega=0.5$.
4.1.2. Baselines
We compare our method with previous subject-driven generation approaches, including zero-shot methods (IP-Adapter(Ye et al., 2023), MS-Diffusion(Wang et al., 2024a), OmniGen(Xiao et al., 2024)) and training-free methods (FreeCustom(Ding et al., 2024), DiptychPrompting(Shin et al., 2024)). Detailed descriptions of these methods are provided in Appendix B.
4.1.3. Datasets and Evaluation
For qualitative evaluation, we use reference images from DreamBench(Ruiz et al., 2023), CustomConcept101(Kumari et al., 2023), and Mix-of-Show(Gu et al., 2023) for subject-driven generation. For quantitative evaluation, we generate 30×25×4=3,000 images using 30 subjects and 25 prompt templates from DreamBench(Ruiz et al., 2023), with 4 random seeds per subject-prompt pair. We evaluate performance from two perspectives: text alignment and image alignment. For text alignment, we report results using CLIP(Radford et al., 2021) and ImageReward(Xu et al., 2024) metrics. For image alignment, we use GroundedSAM(Ren et al., 2024) to extract segments of reference and generated subjects, respectively, and compute similarity scores between these segments using CLIP(Radford et al., 2021) and DINOv2(Oquab et al., 2023).
4.2. Quantitative Results
Image-alignment Text-alignment
Method CLIP-I DINO CLIP-T ImageReward
FreeCustom 0.8308 0.6107 0.3246 0.6223
IP-Adapter 0.8920 0.7696 0.3048 0.7444
MS-Diffusion 0.9023 0.7977 0.3254 1.3405
OmniGen 0.9113 0.8167 0.3256 1.4926
DiptychPrompting 0.8924 0.7971 0.3291 1.5728
Ours $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$
Table 1. Quantitative Comparison of Subject-Driven Generation Methods. Without requiring model fine-tuning or additional training, our FreeGraftor significantly outperforms existing zero-shot and training-free approaches in both image-alignment and text-alignment.
Image-alignment Text-alignment
Variant CLIP-I DINO CLIP-T ImageReward
w/o Init 0.8529 0.7138 0.3236 1.6320
w/o Graft 0.8463 0.6572 0.3306 1.6434
w/o Match 0.9504 0.8974 0.3232 1.5258
w/ replace 0.8476 0.7010 0.3174 1.4551
Ours $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$
Table 2. Quantitative Comparison of Ablation Variants. Absence of structure-consistent initialization or feature grafting leads to degraded image alignment, while omitting semantic matching or replacing patch features results in deteriorated text alignment.
As shown in Table 1, FreeGraftor demonstrates superior image alignment and text alignment compared to prior methods. The significant improvement in image alignment indicates our method effectively preserves the identity of reference subjects: structure-consistent initialization retains geometric priors, while feature grafting successfully transfers fine-grained visual details. The high text alignment stems from leveraging the base text-to-image model without fine-tuning or training auxiliary modules, thereby preserving its original capabilities. Furthermore, FreeGraftor employs a gentle attention fusion strategy to inject subject features, minimizing perturbations to vanilla attention computation.
4.3. Qualitative Results
4.3.1. Single-Subject Generation
Figure 5 presents the generation results for single-subject generation using FreeGraftor and other methods. FreeCustom and IP-Adapter fail to retain fine visual details of the reference subject effectively. While MS-Diffusion, OmniGen, and DiptychPrompting partially recover reference characteristics, they struggle with pixel-level fidelity. FreeGraftor achieves pixel-accurate detail preservation, such as logos on the can, numerals on the clock, and text on the teddy bear. For non-rigid subjects (e.g., dogs, plush toys), FreeGraftor generates natural pose variations aligned with text guidance, demonstrating the flexibility and robustness of our feature grafting technique.
4.3.2. Multi-Subject Generation
As IP-Adapter and DiptychPrompting do not support multi-subject generation, we compare FreeGraftor with FreeCustom, MS-Diffusion, and OmniGen. As shown in Figure 6, FreeCustom exhibits severe attribute confusion (e.g., facial structures of the cat and the dog, hairstyles of the man and the woman). MS-Diffusion and OmniGen show higher fidelity but still lack visual faithfulness for small objects (e.g., the cup). FreeGraftor preserves all reference details while enabling flexible text-guided control, highlighting its effectiveness in multi-subject generation.
4.4. Ablation Study
4.4.1. Component Ablation
To validate the effectiveness of each component in our method, we compare several ablation variants, as shown in Figure 7. (a) When structure-consistent initialization is removed, the generated images inherit the reference subject’s appearance (e.g., colors, patterns) but fail to maintain structural consistency (e.g., distorted body proportions in dogs). (b) Without feature grafting, the generated subjects lack identifiable correspondence with the reference subjects. (c) If semantic matching is omitted and reference keys and values are directly appended during attention computation, the generated subjects exhibit rigid poses and artifacts. (d) Replacing position-constrained attention fusion with direct feature replacement introduces incoherent edges and distortions. Quantitative results in Table 2 confirm that both structure-consistent initialization and feature grafting significantly improve image alignment, as they are critical for preserving subject identity. Skipping semantic matching or positional fusion degrades text alignment due to reduced image quality.
4.4.2. Hyperparameter Analysis
We conduct additional ablation experiments to analyze the impact of similarity threshold $\tau$ and cycle consistency threshold $\delta$ on semantic matching. Appropriate thresholds alleviate mismatches (e.g., unnatural expansion of the backpack), while overly strict thresholds ($\tau\uparrow$ or $\delta\downarrow$) discard critical visual details (e.g., missing badges on the backpack, Figure 8).
We further verify the effectiveness of dropout in mitigating structural overfitting. As shown in Figure 9, moderate dropout enhances pose flexibility while preserving subject identity, whereas excessive dropout intensities ($\omega>0.6$) erode identity retention. These findings demonstrate that through careful hyperparameter selection, our method achieves an optimal balance between pose flexibility and detail preservation.
5. Conclusion
We present FreeGraftor, a novel training-free framework for subject-driven generation that bridges the critical gap between subject fidelity and computational efficiency. By leveraging Semantic-Aware Feature Grafting, FreeGraftor seamlessly injects the appearance of reference subjects into generated scenes. Additionally, our Structure-Consistent Initialization strategy preserves reference geometry and enhances the robustness of semantic matching. Extensive experiments demonstrate that FreeGraftor achieves pixel-level detail preservation and flexible text-guided control without requiring any training or test-time optimization. Our framework can be seamlessly extended to multi-subject generation, faithfully retaining visual details of all reference subjects while incurring no additional computational overhead.
Appendix A Pseudo Code
Input: Reference image $I^{\text{ref}}$, text prompt $P$
Output: Generated image $I^{\text{gen}}$
Stage 1: Collage Construction
1. Template Generation
$I^{\text{tmp}}\leftarrow\text{T2I}(P)$ ;
  // Base text-to-image model
2. Template Subject Processing
$B^{\text{tmp}}\leftarrow\text{GroundingDINO}(I^{\text{tmp}},P)$ ;
  // Bounding box
$M^{\text{tmp}}\leftarrow\text{SAM}(I^{\text{tmp}},B^{\text{tmp}})$ ;
  // Segmentation mask
$I^{\text{erase}}\leftarrow\text{LaMa}(I^{\text{tmp}},M^{\text{tmp}})$ ;
  // Inpainting
3. Reference Subject Extraction
[…]
$M^{\text{ref}}\leftarrow\text{SAM}(I^{\text{ref}},B^{\text{ref}})$
$I^{\text{crop}}\leftarrow\text{CropWithMask}(I^{\text{ref}},M^{\text{ref}})$ ;
  // Mask-based cropping
4. Collage Assembly
$I^{\text{collage}}\leftarrow\text{ResizeAndPaste}(I^{\text{erase}},I^{\text{crop}},B^{\text{tmp}})$ ;
  // Scale & align
Stage 2: Inversion
5. $\epsilon^{\text{init}},\mathcal{F}^{\text{ref}}\leftarrow\text{FireFlow}(I^{\text{collage}})$ ;
  // Noise & feature extraction
Stage 3: Generation
6. $I^{\text{gen}}\leftarrow\text{SAFG}(\epsilon^{\text{init}},\mathcal{F}^{\text{ref}},P)$ ;
  // Feature grafting
return $I^{\text{gen}}$
Algorithm 1 FreeGraftor Pipeline
Input: Reference features $\mathcal{F}^{\text{ref}}$, noise $\epsilon^{\text{init}}$, text prompt $P$
Output: Modified attention features $H_{\text{out}}^{+}$
Semantic Matching
1. For each timestep $t$:
foreach reference patch $i\in M^{\text{ref, pre}}$ do
       foreach generated patch $j$ do
             $s_{ij}\leftarrow\frac{H_{i}^{\text{img,ref}}\cdot H_{j}^{\text{img,gen}}}{\|H_{i}^{\text{img,ref}}\|_{2}\|H_{j}^{\text{img,gen}}\|_{2}}$
             $m(i)\leftarrow\arg\max_{j}s_{ij}$
2. Apply filtering:
$M^{\text{ref}}\leftarrow\text{ThresholdFilter}(s_{ij},\tau)\odot\text{CycleConsistency}(d_{i},\delta)$
Position-Constrained Fusion
3. Concatenate keys/values:
$K^{\text{cat}}\leftarrow[K^{\text{txt}};K^{\text{img,gen}};K_{M}^{\text{img,ref}}]$
$V^{\text{cat}}\leftarrow[V^{\text{txt}};V^{\text{img,gen}};V_{M}^{\text{img,ref}}]$
4. Bind positional embeddings:
$PE^{\text{cat}}\leftarrow\text{Concat}(PE^{\text{txt}},PE^{\text{img,gen}},\{PE_{m(i)}^{\text{img,gen}}\})$
5. Compute revised attention:
$\tilde{Q}\leftarrow PE\odot Q$
$\tilde{K}^{\text{cat}}\leftarrow PE^{\text{cat}}\odot K^{\text{cat}}$
$A^{+}\leftarrow\text{Softmax}(\tilde{Q}(\tilde{K}^{\text{cat}})^{\top}/\sqrt{d})$
$H_{\text{out}}^{+}\leftarrow A^{+}V^{\text{cat}}$
return $H_{\text{out}}^{+}$
Algorithm 2 Semantic-Aware Feature Grafting
Appendix B Baselines
In both qualitative and quantitative comparisons, we only compared our FreeGraftor with training-free approaches and publicly available zero-shot models. Detailed descriptions of these methods are provided below.
  • IP-Adapter(Ye et al., 2023) introduces a lightweight module that projects visual features from an image encoder directly into the cross-attention layers of a pre-trained text-to-image diffusion model. This enables the seamless integration of image prompts with textual conditions without the need for extra fine-tuning.
  • MS-Diffusion(Wang et al., 2024a) tackles multi-subject personalization by using layout guidance and grounding tokens to ensure each subject’s details are preserved and positioned in defined regions. Its specialized cross-attention mechanism effectively minimizes subject conflicts while achieving high visual fidelity in a zero-shot setting.
  • OmniGen(Xiao et al., 2024) unifies diverse image generation tasks, including text-to-image synthesis, editing, subject-driven generation and even classical vision operations—within a single diffusion framework based solely on a VAE and transformer. Trained on a large-scale unified dataset, it avoids additional encoders and transfers knowledge robustly across tasks and domains.
  • FreeCustom(Ding et al., 2024) provides a tuning-free approach for generating multi-concept images by leveraging a multi-reference self-attention mechanism that integrates features from different reference images. With its weighted masking strategy to suppress extraneous details, it efficiently composes high-quality images without further training.
  • DiptychPrompting(Shin et al., 2024) reinterprets subject-driven generation as a text-conditioned inpainting task using a diptych layout, where the left panel holds a reference subject (with background removed) and the right panel is generated from a textual prompt. By enhancing the attention between panels, it captures fine-grained subject details in a zero-shot manner.
Appendix C Additional Experiments
C.1. Efficiency Analysis
Table 3. Efficiency Comparison of Subject-Driven Generation Methods
Method Base Model Time (s) Memory (MiB)
FreeCustom Stable Diffusion v1.5 22.02 14670
IP-Adapter FLUX.1-dev 9.58 40788
MS-Diffusion SDXL $\mathbf{8.89}$ 14654
OmniGen Phi-3 46.13 $\mathbf{12638}$
DiptychPrompting FLUX.1-dev 52.11 46858
Ours FLUX.1-dev 42.62 41106
We compare the efficiency of FreeGraftor with test-time subject-driven generation methods (including zero-shot and training-free approaches) on an NVIDIA A40 GPU. All methods generate 512×512 resolution images under identical conditions, with measurements of peak GPU memory consumption and inference time.
As shown in Table 3, among methods built on the FLUX.1-dev base model, FreeGraftor significantly outperforms DiptychPrompting in both time and memory efficiency, while demonstrating comparable memory usage to IP-Adapter. Across all evaluated methods, OmniGen achieves the lowest memory footprint but requires substantially longer inference time. MS-Diffusion attains the fastest inference speed and second-lowest memory consumption after OmniGen, though these efficiency gains come at the expense of reduced subject fidelity and degraded visual quality. Notably, our FreeGraftor maintains an optimal balance between generation quality and computational efficiency, delivering robust performance without compromising subject faithfulness or visual details.
Method Text-alignment Image-alignment
FreeCustom 1.813 1.638
IP-Adapter 3.090 2.183
MS-Diffusion 3.458 2.905
OmniGen 3.905 3.368
DiptychPrompting 4.633 4.280
Ours $\mathbf{4.815}$ $\mathbf{4.773}$
Table 4. Results of user study. Our method is the most preferred by users, achieving the highest ratings in both text and image alignment.
C.2. User Study
We conduct a user study assessing human preference of text-to-image correspondence and subject identity consistency in generated results. Participants are presented with images synthesized by competing methods alongside their corresponding textual prompts and reference subjects. For text alignment evaluation, participants rate the semantic consistency between generated images and prompts on a 5-point Likert scale (1: minimal correspondence, 5: perfect alignment). For image alignment, they evaluate the visual similarity between generated subjects and reference counterparts using an identical scale, with higher scores indicating stronger identity preservation. The study involves a comprehensive evaluation comprising 40 participants, each assessing 60 randomly sampled images. As detailed in Table 4, our method demonstrated significant superiority over baselines in both metrics, validating its effectiveness in balancing textual faithfulness and subject fidelity.
C.3. More visualization results
Figure 10 presents additional qualitative results comparing our method with the baselines. As shown in these results, our method generates text-aligned novel scenes while maintaining subject fidelity. Our FreeGraftor consistently preserves pixel-level visual details, even for challenging text and patterns, highlighting its effectiveness in subject-driven generation.
Extended single-subject generation results with various prompts are visualized in Figure 11. Our method demonstrates comprehensive prompt adaptability, dynamically modulating subject attributes, poses, and scene configurations to synthesize high-fidelity images that achieve both visual harmony and precise semantic alignment with input textual descriptions.
Figure 10. More qualitative comparisons
Figure 11. Single-subject generation results
Appendix D Limitations
While FreeGraftor achieves robust subject-driven generation in most scenarios, its performance partially depends on the reliability of external models (e.g., GroundingDINO, SAM, LaMa) during the collage construction stage. For instance, rare failures in subject detection, imperfect segmentation masks, or residual artifacts from inpainting could propagate to subsequent stages, occasionally leading to suboptimal initialization. Additionally, the current implementation builds upon the FLUX.1 base model, which demands significant GPU memory (e.g., 40+ GB for 1024×1024 resolution), limiting accessibility for users with constrained hardware. While this choice ensures compatibility with state-of-the-art architectures, adapting our framework to lighter MM-DiT variants or exploring model distillation and quantization techniques could broaden its practical applicability.
Appendix E Social Impacts
The proposed FreeGraftor framework has the potential to democratize personalized content creation by enabling users to synthesize high-fidelity images of specific subjects without requiring extensive training. This capability could empower creators in fields such as digital art, advertising, and education to rapidly prototype visual concepts while preserving intricate details (e.g., logos, textures), thereby lowering barriers to professional-grade content generation. However, the ease of transferring visual identities also raises concerns about misuse, such as generating deceptive imagery for disinformation campaigns or unauthorized replication of copyrighted subjects. Additionally, while our method avoids the computational costs of per-subject fine-tuning, the reliance on large pre-trained diffusion models still necessitates significant energy consumption during inference, highlighting the need for ongoing research into efficient deployment strategies. We advocate for ethical guidelines that balance creative freedom with safeguards against malicious applications, such as watermarking synthesized content or restricting access to sensitive reference data.