### v3

# FreeGraftor: Training-Free Cross-Image Feature Grafting for Subject-Driven Text-to-Image Generation

Abstract.
Subject-driven image generation aims to synthesize novel scenes that faithfully preserve subject identity from reference images while adhering to textual guidance. However, existing methods struggle with a critical trade-off between fidelity and efficiency. Tuning-based approaches rely on time-consuming and resource-intensive, subject-specific optimization, while zero-shot methods often fail to maintain adequate subject consistency. In this work, we propose FreeGraftor, a training-free framework that addresses these limitations through cross-image feature grafting. Specifically, FreeGraftor leverages semantic matching and position-constrained attention fusion to transfer visual details from reference subjects to the generated images. Additionally, our framework introduces a novel noise initialization strategy to preserve the geometry priors of reference subjects, facilitating robust feature matching. Extensive qualitative and quantitative experiments demonstrate that our method enables precise subject identity transfer while maintaining text-aligned scene synthesis. Without requiring model fine-tuning or additional training, FreeGraftor significantly outperforms existing zero-shot and training-free approaches in both subject fidelity and text alignment. Furthermore, our framework can seamlessly extend to multi-subject generation, making it practical for real-world deployment. Our code is available at https://github.com/Nihukat/FreeGraftor.
subject-driven image generation, text-to-image synthesis, training-free framework
$c$s: Computing methodologies Image manipulation
1. Introduction
The democratization of text-to-image generation has ignited widespread interest in personalized content creation, where users seek to synthesize novel scenes containing specific visual concepts—be it a cherished pet, a unique artifact, or a fictional character. While diffusion models(Nichol et al., 2021; Saharia et al., 2022; Ramesh et al., 2022; Rombach et al., 2022; Podell et al., 2023) excel at generating diverse imagery from textual prompts, adapting them to preserve subject identity without incurring costly per-subject optimization remains a critical challenge.
Current subject-driven generation methods(Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023; Li et al., 2023; Ye et al., 2023) face a critical dilemma between fidelity and efficiency. Tuning-based methods (e.g., DreamBooth(Ruiz et al., 2023)) achieve high subject consistency by fine-tuning the model on reference images. However, this subject-specific optimization is resource-intensive, time-consuming, prone to catastrophic forgetting, and struggles with multi-subject composition. Zero-shot methods (e.g., IP-Adapter(Ye et al., 2023)) circumvent fine-tuning by using auxiliary modules (e.g., image encoders or adapters) to extract and inject subject identity information into the generation process. While efficient, these methods often fail to retain fine visual details, particularly for novel concepts, as the limited capacity of the auxiliary modules fundamentally restricts high-fidelity reconstruction.
Figure 1. Cross-image semantic correspondence in FLUX.1. For each specified pixel in the reference image, we identify the most similar pixel in the target image within FLUX.1‘s feature space. These corresponding pixel pairs are then visually connected by colored lines.
Critically, pre-trained text-to-image models inherently possess powerful visual representation capabilities, as evidenced by their ability to synthesize intricate textures and structures directly from text. However, current approaches largely overlook this intrinsic potential: tuning-based methods shift the generation distribution via fine-tuning, while zero-shot methods delegate representation learning to additional modules. Instead, directly leveraging the base model’s own rich feature space for cross-image knowledge transfer presents a promising avenue for bridging the fidelity-efficiency gap. Recent advances in appearance transfer(Epstein et al., 2023; Alaluf et al., 2024; Mou et al., 2023, 2024; Go et al., 2024) support this direction, demonstrating that diffusion models can transplant visual attributes (e.g., style, texture) between images through semantic feature matching. For instance, Eye-for-an-eye (Go et al., 2024) successfully transfers localized appearances by establishing cross-image correspondences within the U-Net’s feature space.
Building upon this concept, we identify a key insight: such cross-image semantic correspondences are not limited to U-Net architectures but also exist robustly within the feature space of modern MM-DiT-based models (Esser et al., 2024) (e.g., FLUX.1 (Labs, 2024)). As demonstrated in Fig. 1, computing pixel-wise similarity (detailed in Sec. 3.2) reveals that semantically corresponding pixels exhibit high similarity in FLUX.1’s feature space.
Figure 2. Subject-driven generation results of our FreeGraftor, achieved without any training or tuning process. These results demonstrate the superior properties of FreeGraftor, including (a) pixel-level detail preservation, (b) flexible text-guided control, and (c) support for multiple reference subjects.
Leveraging this insight, we propose FreeGraftor, a training-free framework that resolves the fidelity-efficiency dilemma in subject-driven generation via cross-image feature grafting. Our approach extracts visual features from reference images using rectified flow inversion (Rout et al., 2024; Wang et al., 2024b; Deng et al., 2024) and injects them into the generation process. Specifically, we introduce a novel cross-image feature grafting mechanism that transfers appearance details through semantic matching and position-constrained attention fusion in the feature space of text-to-image diffusion models. To bind reference patches to their corresponding generated patches, we implement a position embedding replication strategy, enabling semantically matched patches to share positional information during attention computation. Furthermore, we identify that proper noise initialization is crucial for robust semantic matching and structural preservation of reference subjects. Thus, we propose a structure-consistent initialization strategy: first constructing a collage containing the reference subject, then inverting it as the initialization noise for generation. Our method achieves pixel-level detail preservation (e.g., text, patterns) with fast inference and low computational overhead, requiring no additional training or optimization.
Extensive experiments demonstrate that our method outperforms existing approaches in both subject consistency and text alignment while maintaining competitive efficiency. When combining multiple subjects, FreeGraftor preserves individual identities and harmoniously integrates all subjects into a coherent scene without layout conflicts or attribute confusion.
Our contributions are summarized as follows:
  • We introduce FreeGraftor, a plug-and-play framework for subject-driven image generation that achieves pixel-level detail preservation and flexible text-guided control without fine-tuning or training.
  • We develop a cross-image feature grafting technique that transfers visual characteristics of reference subjects to semantically corresponding regions in generated images via semantic matching and position-constrained attention fusion.
  • We propose a novel noise initialization strategy that preserves reference geometry through automated collage construction and rectified flow inversion, seamlessly extending our method to multi-subject scenarios.
2. Related Work
2.1. Text-to-Image Diffusion Models
Text-to-image diffusion models(Ramesh et al., 2022; Saharia et al., 2022; Nichol et al., 2021; Balaji et al., 2022) have achieved remarkable success in generating high-quality and diverse images. Early popular open-source models, such as Stable Diffusion(Rombach et al., 2022) and SDXL(Podell et al., 2023), are based on the U-Net(Ronneberger et al., 2015) architecture and utilize cross-attention between image and text to incorporate textual conditions. In contrast, Diffusion Transformer (DiT)(Peebles and Xie, 2023) replaces the U-Net backbone with a Transformer(Vaswani et al., 2017), demonstrating superior scalability and performance. Recent Multimodal-Diffusion Transformers (MM-DiTs), including Stable Diffusion 3(Esser et al., 2024) and FLUX.1(Labs, 2024), follow this design and project text and image tokens into a unified space to compute joint attention, thereby better integrating complex relationships between text and images. In this work, we implement our method on FLUX.1, where joint attention is computed as:
(1) $Q=\left[Q^{\text{txt}};Q^{\text{img}}\right],\quad K=\left[K^{\text{txt}};K^{\text{img}}\right],\quad V=\left[V^{\text{txt}};V^{\text{img}}\right]$
(2) $\tilde{Q}=PE\odot Q,\quad\tilde{K}=PE\odot K$
(3) $A=\operatorname{softmax}\left(\frac{\tilde{Q}\tilde{K}^{\top}}{\sqrt{d}}\right),\quad H_{\text{out}}=AV$
Here, $[;]$ denotes concatenation, $Q$, $K$, $V$ represent queries, keys, values, $PE$ is position embedding, $d$ denotes the dimension of the input features, $A$ denotes attention scores, and $H_{\text{out}}$ is the output of the attention layer.
2.2. Subject-Driven Generation
Subject-driven generation(Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023; Li et al., 2023; Ye et al., 2023) aims to synthesize novel scenes containing user-specified subjects. Early textual inversion methods(Gal et al., 2022; Dong et al., 2022; Voynov et al., 2023; Alaluf et al., 2023; Liu et al., 2023) learn new concepts by iteratively optimizing embeddings. Mainstream approaches fall into two categories: tuning-based and zero-shot. Tuning-based methods(Ruiz et al., 2023; Kumari et al., 2023; Gu et al., 2023) require per-subject fine-tuning to effectively capture visual details of the reference subject, which is resource-intensive and time-consuming, risks overfitting and catastrophic forgetting, and struggles with multi-subject composition. Zero-shot methods(Gal et al., 2023; Wei et al., 2023; Jia et al., 2023; Shi et al., 2024; Arar et al., 2023; Hua et al., 2023; Li et al., 2023; Ye et al., 2023; Ruiz et al., 2024; Wang et al., 2024a; Wu et al., 2025; Mou et al., 2025) introduce additional learnable modules to process extra image inputs, enabling personalization at inference time. Most of them rely on external encoders to compress images into embeddings, achieving only semantic-level alignment and suffering from insufficient visual consistency, particularly for novel concepts. UNO(Wu et al., 2025) and DreamO(Mou et al., 2025) leverage the VAE within base models to encode reference images for better subject consistency, yet they still struggle to retain fine-grained visual details (e.g., text and patterns). Furthermore, zero-shot methods require large-scale retraining when switching to new base models. Moreover, both categories can degrade the performance of base models, resulting in reduced quality of generated images.
Beyond these categories, several training-free methods(Ding et al., 2024; Shin et al., 2024; Kang et al., 2025) aim to harness the in-context learning potential of pre-trained text-to-image models for plug-and-play subject-driven generation. FreeCustom(Ding et al., 2024) concatenates the reference images and noise in the latent space, and enhances the generated image’s focus on reference subjects by implementing a weighted mask strategy within the self-attention mechanism. DiptychPrompting(Shin et al., 2024) extends this paradigm to MM-DiT-based models via a ControlNet(Zhang et al., 2023), and reframes the subject-driven generation task as image completion on a diptych through meticulous prompt design. LatentUnfold(Kang et al., 2025) dispenses with ControlNet dependencies and evolves the diptych into a grid to enhance subject consistency. These methods avoid resource-demanding training or fine-tuning. However, constrained by the limited in-context learning capacity of base models, relying solely on self-attention struggles to fully reconstruct every detail of the reference subject. Our method identifies cross-image semantic correspondences while binding matched patches, reinforcing contextual connections between reference and generated subjects, thereby ensuring consistency across all components of the subject.
2.3. Appearance Transfer
Appearance transfer methods(Epstein et al., 2023; Alaluf et al., 2024; Mou et al., 2023, 2024; Go et al., 2024) aim to combine the appearance of one image with the structure of another. Techniques like Cross-Image (Alaluf et al., 2024) primarily rely on injecting key-value pairs across images to transfer appearance from reference to target positions in the generated image. Dragon-diffusion(Mou et al., 2023) and DiffEditor(Mou et al., 2024) integrate noise guidance with key-value injection. However, these methods often struggle to preserve structure accurately, as attention scores do not reliably reflect semantic correspondences, which can lead to erroneous matches. Methods like DIFT (Tang et al., 2023) address this by identifying fine-grained correspondences through pixel-pair similarity in U-Net’s feature space. Building on this, Eye-for-an-Eye(Go et al., 2024) achieves appearance transfer by directly replacing features of corresponding pixels. Nevertheless, these methods focus on style transfer rather than identity preservation, which can cause geometric inconsistencies (e.g., shape, limb proportions) between reference and generated subjects. Besides, these U-Net-based approaches are incompatible with modern MM-DiT architectures. In contrast, our work explores semantic matching and fine-grained feature transfer within the MM-DiT architecture while employing a dedicated noise initialization strategy to preserve the geometric structure of reference subjects.
Figure 3. Overview of FreeGraftor. First, we construct a collage based on the given text prompt and the reference image (Stage 1). Next, we invert this collage and record its diffusion trajectory (Stage 2). Finally, using the inverted noise as the initial latent representation, FreeGraftor synthesizes the output image through iterative denoising. During this process, the Semantic-aware Feature Grafting (SAFG) module integrates features from the collage to ensure alignment with the reference subject (Stage 3).
3. Method
3.1. Overview
Our framework comprises three stages: collage construction, inversion, and generation (illustrated in Fig. 3). First, we construct a collage using the given reference image and text prompt. Next, we invert the collage to obtain initialization noise for the generation phase while extracting reference features. Finally, these features are injected into the generation process to synthesize the final image that contains the reference subject within the text-described scene.
Input: Reference image $I^{\text{ref}}$, text prompt $P$
Output: Generated image $I^{\text{gen}}$
Stage 1: Collage Construction
1. Template Generation
$I^{\text{tmp}}\leftarrow\text{T2I}(P)$ ;
// Base text-to-image model
2. Template Subject Processing
$B^{\text{tmp}}\leftarrow\text{GroundingDINO}(I^{\text{tmp}},P)$ ;
// Bounding box
$M^{\text{tmp}}\leftarrow\text{SAM}(I^{\text{tmp}},B^{\text{tmp}})$ ;
// Segmentation mask
$I^{\text{erase}}\leftarrow\text{LaMa}(I^{\text{tmp}},M^{\text{tmp}})$ ;
// Inpainting
3. Reference Subject Extraction
$B^{\text{ref}}\leftarrow\text{GroundingDINO}(I^{\text{ref}},\text{"subject"})$
$M^{\text{ref}}\leftarrow\text{SAM}(I^{\text{ref}},B^{\text{ref}})$
$I^{\text{crop}}\leftarrow\text{CropWithMask}(I^{\text{ref}},M^{\text{ref}})$ ;
// Mask-based cropping
4. Collage Assembly
$I^{\text{collage}}\leftarrow\text{ResizeAndPaste}(I^{\text{erase}},I^{\text{crop}},B^{\text{tmp}})$ ;
// Scale & align
Stage 2: Inversion
5. $\epsilon^{\text{init}},\mathcal{F}^{\text{ref}}\leftarrow\text{FireFlow}(I^{\text{collage}})$ ;
// Noise & feature extraction
Stage 3: Generation
6. $I^{\text{gen}}\leftarrow\text{SAFG}(\epsilon^{\text{init}},\mathcal{F}^{\text{ref}},P)$ ;
// Feature grafting
return $I^{\text{gen}}$
Algorithm 1 FreeGraftor Pipeline
Figure 4. Illustration of the Semantic-aware Feature Grafting (SAFG) Module. For each patch in the reference image, it first establishes semantic correspondence in the generated image via feature matching. The module then (1) concatenates the key and value pairs of the reference patch with those of the corresponding generated patch, while (2) copying and applying the position embedding from the generated patch to the reference patch’s key. This mechanism enables effective positional information sharing between corresponding patches while maintaining semantic alignment.
Input: Reference features $\mathcal{F}^{\text{ref}}$, noise $\epsilon^{\text{init}}$, text prompt $P$
Output: Modified attention features $H_{\text{out}}^{+}$
Semantic Matching
1. For each timestep $t$:
foreach reference patch $i\in M^{\text{ref, pre}}$ do
 foreach generated patch $j$ do
    $s_{ij}\leftarrow\frac{H_{i}^{\text{img,ref}}\cdot H_{j}^{\text{img,gen}}}{\|H_{i}^{\text{img,ref}}\|_{2}\|H_{j}^{\text{img,gen}}\|_{2}}$
    $m(i)\leftarrow\arg\max_{j}s_{ij}$
2. Apply filtering:
$M^{\text{ref}}\leftarrow\text{ThresholdFilter}(s_{ij},\tau)\odot\text{CycleConsistency}(d_{i},\delta)$
Position-Constrained Fusion
3. Concatenate keys/values:
$K^{\text{cat}}\leftarrow[K^{\text{txt}};K^{\text{img,gen}};K_{M}^{\text{img,ref}}]$
$V^{\text{cat}}\leftarrow[V^{\text{txt}};V^{\text{img,gen}};V_{M}^{\text{img,ref}}]$
4. Bind positional embeddings:
$PE^{\text{cat}}\leftarrow\text{Concat}(PE^{\text{txt}},PE^{\text{img,gen}},\{PE_{m(i)}^{\text{img,gen}}\})$
5. Compute revised attention:
$\tilde{Q}\leftarrow PE\odot Q$
$\tilde{K}^{\text{cat}}\leftarrow PE^{\text{cat}}\odot K^{\text{cat}}$
$A^{+}\leftarrow\text{Softmax}(\tilde{Q}(\tilde{K}^{\text{cat}})^{\top}/\sqrt{d})$
$H_{\text{out}}^{+}\leftarrow A^{+}V^{\text{cat}}$
return $H_{\text{out}}^{+}$
Algorithm 2 Semantic-Aware Feature Grafting
3.2. Semantic-Aware Feature Grafting
Most subject-driven generation methods rely on learnable parameters to inject reference features into the generation process. Tuning-based approaches embed subject-specific information into pre-trained models, while zero-shot methods employ additional modules for feature extraction and injection—both incurring significant computational overhead.
Inspired by appearance transfer techniques(Epstein et al., 2023; Alaluf et al., 2024; Mou et al., 2023, 2024; Go et al., 2024), we leverage the inherent capabilities of text-to-image diffusion models to transfer visual characteristics of reference subjects to corresponding regions in generated images without incurring additional training costs. To achieve this, we propose the Semantic-Aware Feature Grafting (SAFG) module, which first establishes semantic correspondences between reference and generated patches via semantic matching, then fuses features through position-constrained attention fusion (Fig. 4).
3.2.1. Semantic Matching
Early appearance transfer methods(Alaluf et al., 2024; Mou et al., 2023, 2024) directly concatenate keys and values of reference and generated images to compute cross-image attention. However, high attention scores do not necessarily indicate semantic correspondences across images, and erroneous matches may lead to structural distortions and visual artifacts. DIFT(Tang et al., 2023) revealed that U-Net-based diffusion models contain rich semantic information in their feature space and established cross-image correspondences via cosine similarity-based semantic matching. We extend this principle to MM-DiT-based diffusion models.
Given a reference image $I^{\text{ref}}$ and a predefined binary mask $M^{\text{ref, pre}}$ indicating the reference subject, we extract attention features using rectified flow inversion(Rout et al., 2024; Wang et al., 2024b; Deng et al., 2024) and match reference $H^{\text{ref}}$ and generated $H^{\text{gen}}$ features at the same timestep and attention layer. For each reference patch $i$ within $M^{\text{ref, pre}}$, we compute cosine similarity with all generated patches $j$:
(4) $s_{ij}=\frac{H_{i}^{\text{img,ref}}\cdot H_{j}^{\text{img,gen}}}{\|H_{i}^{\text{img,ref}}\|_{2}\|H_{j}^{\text{img,gen}}\|_{2}},\quad m(i)=\arg\max_{j}s_{ij}$
where $H_{i}^{\text{img,ref}}$ and $H_{j}^{\text{img,gen}}$ denote features of the $i$-th reference patch and $j$-th generated patch at the attention layer input. To ensure reliable matches, we apply two filtering strategies:
i. Similarity Threshold Filtering. For each reference patch $i$, if its maximum cosine similarity $s_{i,m(i)}$ falls below a predefined threshold $\tau$, the match is discarded. This is represented by the binary mask:
(5) $M_{i}^{\text{ref, sim}}=\begin{cases}1,&s_{i,m(i)}\geq\tau\\ 0,&s_{i,m(i)}<\tau\end{cases}$
ii. Cycle Consistency Filtering. Following DreamMatcher(Nam et al., 2024), for each reference patch $i$ (satisfying $M_{i}^{\text{ref, pre}}=1$), we find the most similar reference patch $k$ to its matched generated patch $m(i)$:
(6) $m^{-1}(m(i))=\arg\max_{k:M_{k}^{\text{ref,pre}}=1}\frac{H_{m(i)}^{\text{img,gen}}\cdot H_{k}^{\text{img,ref}}}{\|H_{m(i)}^{\text{img,gen}}\|_{2}\|H_{k}^{\text{img,ref}}\|_{2}}$
Let $\mathbf{p}_{i}^{\text{ref}}$ and $\mathbf{p}_{m^{-1}(m(i))}^{\text{ref}}$ denote the 2D coordinates of patch $i$ and its reverse-matched counterpart. Matches are rejected if their L2 distance exceeds $\delta$:
(7) $d_{i}=\|\mathbf{p}_{i}^{\text{ref}}-\mathbf{p}_{m^{-1}(m(i))}^{\text{ref}}\|_{2},\quad M_{i}^{\text{ref, consi}}=\begin{cases}1,&d_{i}\leq\delta\\ 0,&d_{i}>\delta\end{cases}$
Combining these filters with the predefined mask yields the final reference image mask:
(8) $M^{\text{ref}}=M^{\text{ref, pre}}\odot M^{\text{ref, sim}}\odot M^{\text{ref, consi}}$
3.2.2. Position-Constrained Attention Fusion
After obtaining the filtered reference mask $M^{\text{ref}}$, we transfer features from the retained reference patches. A naive approach is to directly replace the features of the corresponding patches. However, this may introduce artifacts along object boundaries or even cause image distortion (as shown in Fig. 7). Unlike U-Net, which models positional relationships through convolutional operations, MM-DiT distinguishes patch proximity and orientation via position embeddings (PE). This enables us to align reference patches with their target positions in the generated image by editing position embeddings.
Specifically, we concatenate both the keys and values of the reference patch to those of the generated patch, along with the position embeddings corresponding to the generated patch associated with the reference patch. This design enables the reference patch and its corresponding generated patch to share identical positional information. Let $K^{\text{txt}}$ and $V^{\text{txt}}$ denote the text embeddings’ keys and values, respectively; $K^{\text{img, gen}}$ and $V^{\text{img, gen}}$ represent the generated image’s keys and values; and $K_{M}^{\text{img, ref}}$ and $V_{M}^{\text{img, ref}}$ correspond to the keys and values of the reference patches satisfying $M^{\text{ref}}$. We concatenate these as follows:
(9) $K^{\text{cat}}=[K^{\text{txt}};K^{\text{img,gen}};K_{M}^{\text{img,ref}}],$
$V^{\text{cat}}=[V^{\text{txt}};V^{\text{img,gen}};V_{M}^{\text{img,ref}}]$
For each retained reference patch $i$, we extract the corresponding position embedding $PE_{m(i)}^{\text{img, gen}}$ based on the matching relationship $m(i)$. The new position embedding is then formed by concatenating the following embeddings:
(10) $PE^{\text{cat}}=[PE^{\text{txt}};PE^{\text{img,gen}};\{PE_{m(i)}^{\text{img,gen}}\mid M_{i}^{\text{ref}}=1\}]$
The revised attention computation is formulated as:
(11) $\tilde{Q}=PE\odot Q,\quad\tilde{K}^{\text{cat}}=PE^{\text{cat}}\odot K^{\text{cat}}$
(12) $A^{+}=\text{softmax}\left(\frac{\tilde{Q}(\tilde{K}^{\text{cat}})^{\top}}{\sqrt{d}}\right),\quad H_{\text{out}}^{+}=A^{+}V^{\text{cat}}$
Here, $A^{+}$ represents the updated attention scores, and $H_{\text{out}}^{+}$ denotes the output of the attention layer. By integrating semantic matching and the sharing of position embeddings, our SAFG module effectively binds features between matching patches in the reference and generated images, facilitating cross-image feature grafting.
3.3. Structure-Consistent Initialization
Attention-based feature migration replicates the appearance of reference subjects (e.g., color, texture) but may not preserve structural details (e.g., shape, body proportions). Moreover, Gaussian noise initialization introduces significant randomness, leading to discrepancies in shape and size between generated subjects and those in the reference image, potentially causing semantic matching failures. To maintain structural integrity and enhance the robustness of semantic matching, we propose a structure-consistent initialization strategy. This involves creating a collage containing the reference subject and then employing an inversion technique to derive the initial noise required for the generation phase.
As illustrated in Fig. 3, given a text prompt (e.g., ”a cat sitting on a chair wearing a birthday hat”), we first use a base text-to-image model to generate a template image that faithfully represents the described scene. To preserve the structure of the reference subject, we replace the corresponding subject in the template image with the reference subject. Specifically, a grounding model localizes the target subject in the template, a segmentation model extracts its mask, and an inpainting model removes the subject from the template. The reference subject is then cropped (using the same grounding and segmentation models), resized, and pasted into the erased region to form a collage. This collage replaces the original reference image as the new reference. Finally, we invert this collage to obtain the initial noise for the generation phase and record the diffusion trajectory during inversion to facilitate feature grafting.
To prevent the generated subject from overfitting the pose of the reference subject, we introduce a dynamic feature dropout strategy. Since structural layouts are predominantly determined during early diffusion steps(Balaji et al., 2022; Alaluf et al., 2024), we reduce the injection of reference features during the early steps while retaining them in the later steps to preserve visual details. At timestep $t$, for each reference patch, we apply dropout with probability $p=\omega t$, where $\omega\in[0,1]$ controls dropout intensity. The dropout mask is represented by a Bernoulli distribution:
(13) $M^{\text{drop}}_{i,j}\sim\text{Bernoulli}(1-\omega t)\quad\forall i,j$
The final reference mask is computed by applying dropout to the reference image mask obtained from semantic matching (Sec. 3.2):
(14) ${M^{\text{ref}}}^{\prime}=M^{\text{ref}}\odot M^{\text{drop}}$
This strategy preserves the reference subject’s structure while aligning it with the text-driven layout and maintaining pose flexibility. For multi-subject generation, we iteratively replace subjects in the template and invert the multi-subject collage. This avoids simultaneous processing of multiple reference images and eliminates additional computational overhead.
4. Experiments
4.1. Experimental Setup
4.1.1. Implementation Details
Our method is implemented on FLUX.1-dev. For collage construction, we use Grounding DINO(Liu et al., 2024) to localize subjects, SAM(Kirillov et al., 2023) to segment subjects, and LaMa(Suvorov et al., 2022) to erase the original subject in the template image. FireFlow(Deng et al., 2024) is employed for rectified flow inversion. Template generation, inversion, and final generation each take 25 diffusion steps. During semantic matching, we set the similarity threshold $\tau=0.2$, cycle consistency threshold $\delta=1.5$, and dropout intensity $\omega=0.5$. All experiments are conducted on an NVIDIA L20 GPU.
4.1.2. Baselines
We compare our method with zero-shot methods (BLIP-Diffusion, IP-Adapter(Ye et al., 2023), MS-Diffusion(Wang et al., 2024a), UNO(Wu et al., 2025), DreamO(Mou et al., 2025)) and training-free methods (FreeCustom(Ding et al., 2024), DiptychPrompting(Shin et al., 2024), LatentUnfold(Kang et al., 2025)). Detailed descriptions of these methods are provided below.
  • FreeCustom(Ding et al., 2024) provides a tuning-free approach for generating multi-concept images by leveraging a multi-reference self-attention mechanism along with a weighted mask strategy so that the generated image attends to each reference concept. With its weighted masking strategy to suppress extraneous details. This design efficiently preserves the semantic content of each input concept within context without further training.
  • BLIP-Diffusion(Li et al., 2023) trains a multimodal encoder to produce a unified text-aligned visual representation of the target subject. During training, it learns to leverage this subject representation so the diffusion model can generate new renditions of the subject from text prompts.
  • MS-Diffusion(Wang et al., 2024a) injects reference images and explicit grounding tokens using a CLIP-based feature resampler to preserve fine details for each reference subject. Using the provided layout as guidance, it applies a multi-subject cross-attention mechanism so that each subject’s conditioning affects a specific region of the image, which harmonizes the composition of all subjects while respecting the text prompt.
  • IP-Adapter(Ye et al., 2023) introduces a lightweight module that projects visual features from a CLIP image encoder directly into the cross-attention layers of a pre-trained text-to-image diffusion model. This enables the seamless integration of image prompts with textual conditions without the need for extra fine-tuning.
  • UNO(Wu et al., 2025) first uses the inherent in-context generation ability of MM-DiTs to synthesize large-scale, high-consistency multi-subject paired data. Then it trains a “subject-to-image” model (conditioned on multiple images) with progressive cross-modal alignment and a universal rotary positional embedding. This pipeline yields a model that maintains high consistency while supporting control in both single- and multi-subject generation tasks.
  • DreamO(Mou et al., 2025) is a unified image customization framework built on a MM-DiT. DreamO is trained on a large mixed dataset covering diverse tasks (identity, subject, style, background, etc.), and uses techniques to integrate multiple conditions. It employs a feature-routing constraint that queries the relevant information from each reference, and a placeholder strategy that assigns each condition a dedicated position in the input. DreamO can perform many customization tasks with high quality and flexibly combine different types of control conditions in one model.
  • DiptychPrompting(Shin et al., 2024) reinterprets subject-driven generation as a text-conditioned inpainting task using a diptych layout, where the left panel holds a reference subject (with background removed) and the right panel is generated from a textual prompt. By enhancing the cross-attention between the two halves, DiptychPrompting precisely aligns the generated subject with the reference.
  • LatentUnfold(Kang et al., 2025) reframes subject generation as grid-based image completion: the reference images are simply tiled in a mosaic layout to “activate” the model’s latent identity features. This simple “free lunch” approach is enhanced with a cascade attention design and a meta-prompting technique to boost fidelity.
4.1.3. Datasets and Evaluation
For qualitative evaluation, we use reference images from DreamBench(Ruiz et al., 2023), Mix-of-Show(Gu et al., 2023) and CustomConcept101(Kumari et al., 2023) for subject-driven generation. For quantitative evaluation, we generate 30×25×4=3,000 images using 30 subjects and 25 prompt templates from DreamBench(Ruiz et al., 2023), with 4 random seeds per subject-prompt pair. We evaluate performance from two perspectives: text alignment and image alignment. For text alignment, we report results using CLIP(Radford et al., 2021) and ImageReward(Xu et al., 2024) metrics. For image alignment, we use GroundedSAM(Ren et al., 2024) to extract segments of reference and generated subjects, respectively, and compute similarity scores between these segments using CLIP(Radford et al., 2021) and DINOv2(Oquab et al., 2023).
4.2. Quantitative Results
As shown in Table 1, methods using smaller base models (FreeCustom and BLIP-Diffusion) exhibit limitations in both image and text alignments. MS-Diffusion and IP-Adapter achieve lower image alignment than UNO and DreamO, suggesting that reliance solely on CLIP-based external encoders is insufficient for preserving intricate subject details. Training-free methods (DiptychPrompting and LatentUnfold) underperform in image alignment compared to trained zero-shot alternatives (UNO and DreamO), which indicates that the inherent context learning capacity of the base model itself becomes a limiting factor. Crucially, our method achieves the highest alignment across all metrics, demonstrating superior subject fidelity and textual adherence, validating its effectiveness for subject-driven generation.
Image-alignment Text-alignment Efficiency
Method Base Model CLIP-I DINO CLIP-T ImageReward Time (s) Memory (MiB)
FreeCustom SD v1.5 0.8308 0.6107 0.3246 0.6223 14.93 12068.73
BLIP-Diffusion SD v1.5 0.8621 0.7364 0.2949 0.2360 1.44 3641.84
MS-Diffusion SDXL 0.9023 0.7977 0.3254 1.3405 4.02 12268.41
IP-Adapter FLUX.1-dev 0.8920 0.7696 0.3048 0.7444 7.45 38690.51
UNO FLUX.1-dev 0.9069 0.8237 0.3151 1.2788 14.02 35248.43
DreamO FLUX.1-dev 0.9159 0.8400 0.3217 1.6234 12.15 34431.61
DiptychPrompting FLUX.1-dev 0.8924 0.7971 0.3291 1.5728 43.27 44928.55
LatentUnfold FLUX.1-dev 0.8741 0.7755 0.3170 1.3929 104.90 35402.22
Ours FLUX.1-dev $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$ 40.93 28993.33
Table 1. Quantitative Comparison of Subject-Driven Generation Methods. Without requiring model fine-tuning or additional training, our FreeGraftor significantly outperforms existing zero-shot and training-free approaches in both image-alignment and text-alignment.
In additon, we compare the efficiency of FreeGraftor with test-time subject-driven generation methods (including zero-shot and training-free approaches) on an NVIDIA L20 GPU. All methods generate 512×512 resolution images under identical conditions, with measurements of peak GPU memory consumption and inference time. As shown in Table 1, among methods built on the FLUX.1-dev base model, FreeGraftor significantly outperforms all other approaches in terms of memory usage, while also achieving shorter inference time compared to the other two training-free methods (DiptychPrompting and LatentUnfold). Compared to zero-shot methods, our approach is also competitive. While IP-Adapter, UNO, and DreamO have faster inference speeds than FreeGraftor, they consume more GPU memory. Due to their smaller parameter sizes, FreeCustom, BLIP Diffusion, and MS-Diffusion achieve high efficiency in both time and memory, but this comes at the cost of reduced subject consistency and lower visual quality. Notably, our FreeGraftor maintains an optimal balance between generation quality and computational efficiency, delivering robust performance without compromising subject faithfulness or visual details.
Figure 5. Generation results with single reference subject using different methods. Our FreeGraftor achieves pixel-level detail preservation (e.g., text and patterns) while allowing flexible text-guided control (e.g., poses of teddy bears and dogs).
Figure 6. Generated results with different text prompts under the same reference image.
Figure 7. Generation results with multiple reference subjects using different methods. Our FreeGraftor preserves the visual details of all reference subjects while producing high-quality images well-aligned with the text.
4.3. Qualitative Results
4.3.1. Single-Subject Generation
Fig. 5 presents the results for single-subject generation using FreeGraftor and other methods. FreeCustom, BLIP-Diffusion and IP-Adapter fail to retain fine visual details of the reference subject effectively. While other baselines partially recover reference characteristics, they struggle with fine-grained fidelity. FreeGraftor achieves pixel-level detail preservation, such as logos on the can, numbers on the clock, and text on the teddy bear. For non-rigid subjects (e.g., dogs, plush toys), FreeGraftor generates natural pose variations aligned with text guidance, demonstrating the flexibility and robustness of our feature grafting technique.
Extended single-subject generation results with various prompts are visualized in Fig. 6. Our method demonstrates comprehensive prompt adaptability, dynamically modulating subject attributes, poses, and scene configurations to synthesize high-fidelity images that achieve both visual harmony and precise semantic alignment with input textual descriptions.
4.3.2. Multi-Subject Generation
As shown in Fig. 7, FreeCustom and MS-Diffusion struggle to achieve sufficient subject consistency. While UNO and DreamO demonstrate higher fidelity, they suffer from subject omission (e.g., the barn and the table) and attribute confusion (e.g., confusion between the cat and the dog). FreeGraftor preserves all reference details while enabling flexible text-guided control, highlighting its effectiveness in multi-subject generation.
Figure 8. Generation results of different ablation variants. (a) Without structure-consistent initialization. (b) Without feature grafting. (c) Using direct key-value concatenation in attention layers instead of semantic matching. (d) Using direct patch feature replacement instead of position-constrained attention fusion. (e) Our full method.
Image-alignment Text-alignment
Variant CLIP-I DINO CLIP-T ImageReward
w/o Structure-consistent Initialization 0.8529 0.7138 0.3236 1.6320
w/o Feature Grafting 0.8463 0.6572 0.3306 1.6434
w/o Semantic Matching 0.9504 0.8974 0.3232 1.5258
w/ Feature Replacement 0.8476 0.7010 0.3174 1.4551
Ours $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$
Table 2. Quantitative Comparison of Ablation Variants. Absence of structure-consistent initialization or feature grafting leads to degraded image alignment, while omitting semantic matching or replacing patch features results in deteriorated text alignment.
4.4. Ablation Study
4.4.1. Component Ablation
To validate the effectiveness of each component in our method, we compare several ablation variants, as shown in Fig. 8. (a) When structure-consistent initialization is removed, the generated images inherit the reference subject’s appearance (e.g., colors and patterns) but fail to maintain structural consistency (e.g., distorted body proportions in dogs). (b) Without feature grafting, there is no visual correlation between the generated and reference subjects. (c) If semantic matching is omitted and reference keys/values are directly appended during attention computation, the generated subjects exhibit rigid poses and artifacts. (d) Replacing position-constrained attention fusion with direct feature replacement introduces incoherent edges and distortions. Quantitative results in Table 2 confirm that both structure-consistent initialization and feature grafting significantly improve image alignment, as both are critical for preserving subject identity. Skipping semantic matching or positional fusion degrades text alignment due to reduced image quality.
Figure 9. Generation results under different similarity threshold ($\tau$) and cycle-consistency threshold ($\delta$) settings. Higher $\tau$ and lower $\delta$ lead to more stringent filtering.:
Figure 10. Generation results under different dropout intensity ($\omega$) settings. Higher $\omega$ improves text alignment but degrades subject consistency.:
4.4.2. Hyperparameter Analysis
We conduct additional ablation experiments to analyze the impact of similarity threshold $\tau$ and cycle consistency threshold $\delta$ on semantic matching. Appropriate thresholds alleviate mismatches (e.g., unnatural expansion of the backpack), while overly strict thresholds ($\tau\uparrow$ or $\delta\downarrow$) discard critical visual details (e.g., missing badges on the backpack, Fig. 9).
We further verify the effectiveness of dropout in mitigating structural overfitting. As shown in Fig. 10, moderate dropout enhances pose flexibility while preserving subject identity, whereas excessive dropout intensities ($\omega>0.6$) erode identity retention. These findings highlight the necessity of balancing pose flexibility and detail preservation through parameter tuning.
Figure 11. Comparison between DreamBooth and our method on FLUX.1. Without computation-intensive fine-tuning, our method achieves performance comparable to DreamBooth and even surpasses it in capturing visual details such as text and patterns.
Figure 12. Generation results of (a) Stable Diffusion (b)Stable Diffusion with DreamMatcher (c) DreamBooth on Stable Diffusion (d) DreamBooth on Stable Diffusion with DreamMatcher (e) Ours.
4.5. Comparison with DreamBooth
We compare FreeGraftor with DreamBooth, a widely-used tuning-based method. Due to limitations in computational resources, we implemented DreamBooth using LoRA instead of full parameter fine-tuning on FLUX.1-Dev, with a LoRA rank of 16.
As shown in Fig. 11, although fine-tuned for several tens of minutes on the reference images of each subject, DreamBooth still struggles to fully reproduce the visual details of the reference subject (e.g., text and patterns). Additionally, fine-tuning on limited samples can lead to overfitting of the background, compromising its ability to follow textual semantics. In contrast, without resource-intensive fine-tuning, our method achieves pixel-level detail preservation and faithful text alignment.
Image-alignment Text-alignment
Variant CLIP-I DINO CLIP-T ImageReward
DreamBooth 0.9261 0.8776 0.3123 1.3232
Ours $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$
Table 3. Quantitative Comparison of DreamBooth and our method on FLUX.1-Dev.
Furthermore, we quantitatively compared FreeGraftor with DreamBooth using the same experimental setup (Sec. 4.1.3). Table 3 shows our approach delivers significant performance advantages beyond just efficiency gains. FreeGraftor resolves the fidelity-efficiency dilemma in subject-driven generation, going beyond merely alleviating the limitations of training-free methods.
4.6. Comparison with DreamMatcher
The most comparable approach to ours is likely DreamMatcher(Nam et al., 2024), yet the two remain fundamentally distinct. Firstly, DreamMatcher is not a standalone subject-driven generation method but rather a plugin. It requires integration with tuning-based approaches (e.g., DreamBooth), which involve time-consuming and resource-intensive test-time optimization. In contrast, our method is a standalone subject-driven generation pipeline without any fine-tuning or training. Secondly, DreamMatcher adopts a direct attention feature replacement mechanism for identity injection, which tends to introduce artifacts along subject boundaries. In contrast, our approach leverages cross-image attention with shared spatial information, employing a more refined methodology that effectively mitigates image distortion.
Since DreamMatcher is tailored specifically for Stable Diffusion 1.5 and cannot be directly migrated to other network architectures, we implemented DreamMatcher on Stable Diffusion 1.5 following the official code. As show in Fig. 12, when employed in conjunction with DreamBooth, DreamMatcher effectively rectifies erroneous attribute binding(e.g., the yellow number ”3” on the clock), thereby enhancing the fidelity of reconstructed visual details. However, when applied directly to Stable Diffusion, DreamMatcher is unable to fully reconstruct the subject from the reference image, owing to a lack of prior structural knowledge of the target subject. Additionally, some images generated by DreamMatcher exhibited noticeable distortions (e.g., green and blue artifacts around the backpack).
Image-alignment Text-alignment
Variant CLIP-I DINO CLIP-T ImageReward
Stable Diffusion + DreamMatcher 0.7830 0.5336 0.3232 0.6387
DreamBooth + DreamMatcher 0.8912 0.7918 0.3021 0.1767
Ours $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$
Table 4. Quantitative Comparison of DreamMatcher and our method.
Quantitative comparisons under identical settings (Sec. 4.1.3) confirm our superiority, as shown in Table 4. Our method achieves significant performance gains through two key designs. First, the structure-consistent noise initialization produces reasonable image layouts while faithfully preserving reference geometry, which ensures robust semantic matching and detail retention. Second, our position-constrained attention fusion mechanism enables harmonious integration of features, resulting in more natural-looking outputs. Ablation studies (Sec. 4.4.1) confirm the effectiveness of these core components.
5. Conclusion
We present FreeGraftor, a novel training-free framework for subject-driven generation that bridges the critical gap between subject fidelity and computational efficiency. By leveraging Semantic-Aware Feature Grafting, FreeGraftor seamlessly transfers the appearance of reference subjects into generated scenes. Additionally, our Structure-Consistent Initialization strategy preserves reference geometry and enhances the robustness of semantic matching. Extensive experiments demonstrate that FreeGraftor achieves pixel-level detail preservation and flexible text-guided control while requiring no training or test-time optimization. Our framework can be seamlessly extended to multi-subject generation, faithfully retaining visual details of all reference subjects while incurring no additional computational overhead.
Acknowledgements.
This work was supported by Beijing Natural Science Foundation (L247010), NSFC (No.61906018).